# GM3: A General Physical Model for Micro-Mobility Vehicles

### v1

@@ Abstract
Modeling the dynamics of micro-mobility vehicles (MMV) is becoming increasingly important for training autonomous vehicle systems and building urban traffic simulations. However, mainstream tools rely on variants of the Kinematic Bicycle Model (KBM) [28, 16, 24, 9] or mode-specific physics that miss tire slip, load transfer, and rider/vehicle lean. To our knowledge, no unified, physics-based model captures these dynamics across the full range of common MMVs and wheel layouts. We propose the “Generalized Micro-mobility Model” (GM3), a tire-level formulation based on the tire brush representation [1, 2, 3] that supports arbitrary wheel configurations, including single/double track and multi-wheel platforms. We introduce an interactive model-agnostic simulation framework that decouples vehicle/layout specification from dynamics to compare the GM3 with the KBM and other models, consisting of fixed step RK4 integration, human-in-the-loop and scripted control, real-time trajectory traces and logging for analysis. We also empirically validate the GM3 on the Stanford Drone Dataset’s deathCircle (roundabout) scene [8] for biker, skater, and cart classes.
@@ I INTRODUCTION
Micro-mobility vehicles (MMVs) such as bicycles and scooters are becoming increasingly central to today’s urban traffic. In 2023, people took 133 million trips on shared micro-mobility across the US and 24 million trips in Canada, led by the rise of dockless and station-based shared systems [25]. Beyond bikes and e-scooters, adoption of other forms of micro-mobility, including skateboards, and low-speed vehicles (LSVs), are also on the rise [27, 15].
@@
Furthermore, many state-of-the-art systems default to some variant of the Kinematic Bicycle Model (KBM) [28, 16, 24, 9] due to its simplicity and low computational cost, performing well under low-speed, low-curvature conditions. However, the KBM yields unrealistic trajectories when MMVs perform aggressive or high-curvature maneuvers, such as sharp turns, hard braking, and rapid lane changes. The model does not capture tire slip, load transfer, or rider lean, and its single-track geometry is not generalizable to other forms of micro-mobility with alternate wheel configurations. These limitations, therefore, make the KBM insufficient for simulating the full range of MMV behaviors that autonomous systems must anticipate in mixed traffic.
@@
  • An interactive simulation framework that implements the GM3. The framework enables plug-and-play specification of MMV dynamics models, vehicle geometries, and parameters (e.g. vehicle, tire, lean) and provides real-time control and trajectory visualization.
@@ II RELATED WORKS
Kinematic Bicycle Model (KBM) is a simple and computationally inexpensive model that is accurate for low speeds [4]. However, it has considerable limitations in modeling high-speed dynamics and trajectories with high curvature. [7] demonstrated that KBM only yielded consistent results for lateral acceleration $a_{y}<0.5\mu g$. In an experiment that aimed to improve KBM for higher lateral accelerations with an additional state variable [10], the results show vehicle heading error growing significantly with the curvature due to under-steering in the simulation. This is because KBM does not account for lean or rider dynamics. KBM is also not generalizable to other MMVs with different wheel configurations.
Using a Tire Model as a basis, vehicles with varying numbers of wheels in different positions can be simulated by applying the same tire model with adjusted parameters for the tires’ physical properties and transforming the forces obtained from the brush model into the vehicle body frame. [6] performed validation experiments of the brush model with a Chrysler Voyager, and the trajectories from the experiments closely matched the corresponding simulation results for low speeds. In [11], the brush model and a thermo-physical brush model were validated on a Formula SAE vehicle in extreme scenarios, such as high-speed lane changes. The qualitative and quantitative results show agreement between both models and the experimental data, with an average fitting error of 6.19% for the basic brush model and 6.09% for the thermo-physical model. The semi-empirical Magic Formula – widely used for commercial purposes – was also evaluated against the data and only provided a 0.46% improvement over the brush model. These results demonstrate the viability of the brush model for simulating vehicle dynamics at low and high speeds.
Several implementations of the brush model and other tire models have been designed for specific vehicles. Project Chrono [12] provides a Vehicle class with templates for subsystems, including suspension, steering, and braking, and options for three different classes of tire models: rigid, semi-empirical, and finite element. The user can define a vehicle as a list of axles, each consisting of a set of tires with their own parameters. While this library can handle various wheel configurations, its purpose is to model complex military vehicles and as such, it does not address rider dynamics, and the computational overhead is excessive for MMVs. There is no general model for MMVs that allows the user to customize the wheel layout similarly to Project Chrono, and this paper intends to fill that gap.
@@ III-A Tire Brush Model
The brush model is derived from calculating the deflection of each bristle and integrating over the contact patch between the tire and the road surface, assuming a parabolic pressure distribution for simplicity. Tables I and II include the minimum parameters and variables required to calculate slip forces for a single tire.
@@
$a$ Half contact length m
@@
The brush model uses a 3D coordinate system with $x$ pointing forward along the tire’s trajectory (longitudinal), $y$ pointing right perpendicular to $x$, and $z$ pointing up. For simplicity, we assume the isotropic model where the tire and road properties are the same in every direction ($c_{p}=c_{px}=c_{py}$ and $\mu=\mu_{x}=\mu_{y}$) and we assume a flat road surface. Then, we define the theoretical slip vector $\mathbf{\sigma}$ with practical slip quantities $\kappa=-\frac{V_{sx}}{V_{x}}$ and $\tan\alpha=-\frac{V_{sy}}{V_{x}}$ (where $V_{x}$ and $V_{y}$ are the linear longitudinal and lateral velocities and $V_{sx}$ and $V_{sy}$ are the linear longitudinal and lateral velocities slip velocities).
$\mathbf{\sigma}=-\frac{\mathbf{V_{s}}}{V_{r}}=\begin{pmatrix}\sigma_{x}\\ \sigma_{y}\end{pmatrix}\\ =\begin{pmatrix}\frac{\kappa}{1+\kappa}\\ \frac{\tan\alpha}{1+\kappa}\end{pmatrix}$ (1)
We introduce spin $\varphi$ which represents the turn slip that occurs when a tire follows a curved path. For a tire with camber angle $\gamma$, effective rolling radius $R_{e}$, turning radius $r_{turn}$, and reduction factor $\varepsilon_{\gamma}$ (close to 0 for small tires and close to 1 for large tires), the turn slip is
$\varphi=-\frac{1}{r_{turn}}+(1-\varepsilon_{\gamma})\frac{\sin(\gamma)}{r_{turn}}.$ (2)
@@
$F_{x}$ $=\begin{cases}\mu F_{z}\frac{\sigma_{x}}{\sigma}[3\theta\sigma-3(\theta\sigma)^{2}+(\theta\sigma)^{3}],&\sigma\leq\frac{1}{\theta}\\ \mu F_{z},&\sigma>\frac{1}{\theta}\end{cases}$ (4)
$F_{y}$ $=\begin{cases}\mu F_{z}[3\theta^{*}\sigma_{y}-3(\theta^{*}\sigma_{y})^{2}+(\theta^{*}\sigma_{y})^{3}]+\frac{2}{3}c_{p}a^{3}\varphi,&|\sigma_{y}|\leq\frac{1}{\theta^{*}}\\ \mu F_{z}\text{sgn}(\alpha),&|\sigma_{y}|>\frac{1}{\theta^{*}}\end{cases}$ (5)
$M_{z}$ $=\begin{cases}-\mu F_{z}a\theta^{*}\sigma_{y}[1-3|\theta^{*}\sigma_{y}|+3(\theta^{*}\sigma_{y})^{2}-|\theta^{*}\sigma_{y}|^{3}],&|\sigma_{y}|\leq\frac{1}{\theta^{*}}\\ 0,&|\sigma_{y}|>\frac{1}{\theta^{*}}\end{cases}$ (6)
The values $\frac{1}{\theta}$ and $\frac{1}{\theta^{*}}$ represent the boundary between the adhesion region ($\sigma\leq\frac{1}{\theta}$), where the bristles deform without sliding, and the sliding region ($\sigma>\frac{1}{\theta}$), where the tire has reached its maximum available friction. Here, sgn is the sign function, $\sigma=\sqrt{\sigma_{x}^{2}+\sigma_{y}^{2}}$ and $\theta^{*}$ is the modified tire parameter for camber angle:
$\theta^{*}=\frac{\theta}{1-a\varphi\theta\text{sgn}(\alpha)}$ (7)
@@ IV METHODOLOGY
Figure 2: The GM3 State gets updated with the following steps: (1) load transfer and leaning angle to steer angle conversion for platforms, (2) control assignment for each tire, (3) individual tire processing with the brush model, and (4) force integration and rider lean force application.
@@ IV-B Load Transfer
$F_{z,front}$ $=\frac{mgl_{r}}{n_{f}L}$ (15)
$F_{z,rear}$ $=\frac{mgl_{f}}{n_{r}L}$ (16)
@@
$T_{long}$ $=\frac{ma_{x}h_{cg}}{L}$ (17)
$T_{lat}$ $=\frac{ma_{y}h_{cg}}{W}$ (18)
@@ IV-D Skateboard Truck Geometry
$\delta_{f}$ $=k\phi\sin{\beta}$ (23)
$\delta_{r}$ $=-k\phi\sin{\beta}$ (24)
@@ IV-E Generalization to Diverse Vehicle Platforms
Our GM3 is widely applicable to model and simulate a diverse range of micro-mobility vehicles across arbitrary layouts. Some common modes are shown in Fig. 6, including bicycle, scooter, skateboard, cart/LSV, hoverboard, delta 3-wheeler, tadpole 3-wheeler, 5-wheeled circular platform, and drawn carriage.
((a)) Bicycle // 2-wheel Scooter
@@
((f)) Delta 3-Wheeler
((g)) Tadpole 3-Wheeler
((h)) 5-Wheeled Circular Platform
@@
IV-F Interactive Simulation Framework
To evaluate the proposed GM3 and facilitate comparison with the KBM baseline, we developed an interactive, model-agnostic simulation framework. It is designed to be both a tool for validating the dynamics of different micro-mobility modes and a visualization environment for comparing trajectories produced by different models. Additionally, while this work primarily focuses on comparing GM3 against KBM, the framework is designed to be extensible so additional models can be introduced by adhering to the same interface for control mappings, state, and visual updates.
@@ IV-F Interactive Simulation Framework
At runtime, the simulator advances dynamics with fixed-step fourth-order Runge-Kutta (RK4) integration. At each time step, inputs or controls are read (from keyboard or a script), transformed into model-specific commands, and fed to the appropriate model to update the state. The new state is then used to append time-aligned state and control values to logs and update the visualization which, beyond vehicle effects, extends an accumulating trajectory trace. The trace can be cleared or reset to rerun maneuvers with different configurations. Run histories can be exported to CSV with model-specific fields for analysis.
@@ V EXPERIMENTS
In our experiments, we evaluate the accuracy and correctness of GM3 in reproducing realistic, physically-ground micro-mobility trajectories, as compared to the KBM.
V-A Datasets
@@
Experiments are conducted using the Stanford Drone Dataset (SDD) [8], a large collection of aerial videos recorded over the Stanford campus. Each video is accompanied by frame-level annotations in the form of axis-aligned bounding boxes and categorical labels for eight classes of road users, including pedestrians, cyclists, skaters, carts, cars, and buses. For this work specifically, we focus on the deathCircle (roundabout) scene since it contains diverse micro-mobility interactions as bicyclists, skateboarders, and carts navigate a shared space alongside pedestrians taking high curvature routes. Image-plane trajectories are converted to metric world coordinates using the scene’s pixel-to-meter scale factors [13]. For each annotation, we extract the bottom-midpoint of the bounding box, then resample at the native 30 fps [14] to obtain a uniform step size $\Delta t\approx 0.03$s.
V-B Vehicle Instantiation
@@
Vehicles, one per mode, are instantiated using representative, publicly reported parameters (e.g., wheelbase, track width, tire radius, mass) taken from widely available bicycle, skateboard, and cart specifications on company websites. The bicycle MMV agent uses a single track geometry with lean, the skater uses a skateboard configuration with rider lean, and the cart uses a four-wheel layout and no rider lean. All layouts use load transfer. To prevent overfitting, parameters are fixed per mode and not tuned per track.
V-C Metrics
@@
For each trajectory, we quantify model performance with two measures: (1) Average Displacement Error (ADE) and (2) Discrete Fréchet Distance (DFD). Average Displacement Error (ADE) is computed as the mean Euclidean Distance between the estimated model positions and the corresponding ground truth positions across all time steps. ADE captures how well the simulated trajectory remains aligned with the true path when both are driven by the same underlying control sequence. To account for overall path similarity, we also report the discrete Fréchet distance (DFD).
V-D Results
@@
Table III summarizes ADE and DFD errors across all three modes. GM3 reduces ADE relative to KBM across all modes (Biker: -5.44%, Skater: -9.62%, Cart: -7.32%), which indicates better step-by-step alignment. For DFD, GM3 improves on the Skater (-6.49%) and Cart (-2.4%) modes, but is worse on Biker (+11.56%). Qualitatively, GM3 reproduces better short-term responses to controls due to the tire-level formulation which aligns with the ADE observations. Since Fréchet distance emphasizes global path similarity, the larger lean angles of cyclists relative to skaters and golf carts introduce lateral biases that accumulate along the extracted bottom-midpoint trajectories, which can inflate the distance. However, aggregated across all modes, GM3 performs only around 1% worse than KBM with respect to DFD. Furthermore, the campus setting primarily captures low-speed, cooperative interactions on narrow walkways and roadways (even at the roundabout), so hard braking, rapid lane changes, and sustained high-speed, high-curvature turns tend to be rare. As a result, the reported differences likely understate the advantage GM3 provides in higher-speed scenarios with larger slip angles (e.g., bike lanes, downhill stretches, large evasive maneuvers). Please refer to the supplementary video for live side-by-side comparisons of GM3 vs. KBM for different modes.

### v2

@@ Abstract
Modeling the dynamics of micro-mobility vehicles (MMV) is becoming increasingly important for training autonomous vehicle systems and building urban traffic simulations. However, mainstream tools rely on variants of the Kinematic Bicycle Model (KBM) [13, 4, 12, 11] or mode-specific physics that miss tire slip, load transfer, and rider/vehicle lean. To our knowledge, no unified, physics-based model captures these dynamics across the full range of common MMVs and wheel layouts. We propose the “Generalized Micro-mobility Model” (GM3), a tire-level formulation based on the tire brush representation [svendenius2003brush, deur20043d, svendenius2004brush] that supports arbitrary wheel configurations, including single/double track and multi-wheel platforms. We introduce an interactive model-agnostic evaluation and visualization framework that decouples vehicle/layout specification from dynamics to compare the GM3 with the KBM and other models, consisting of fixed step RK4 integration, human-in-the-loop and scripted control, real-time trajectory traces, and logging for analysis. We also empirically validate the GM3 on the Stanford Drone Dataset’s deathCircle (roundabout) scene [20] for biker, skater, and cart classes.
@@ I INTRODUCTION
Micro-mobility vehicles (MMVs) such as bicycles and scooters are becoming increasingly central to today’s urban traffic. In 2023, people took 133 million trips on shared micro-mobility across the US and 24 million trips in Canada, led by the rise of dockless and station-based shared systems [15]. Beyond bikes and e-scooters, adoption of other forms of micro-mobility, including skateboards and low-speed vehicles (LSVs), is also on the rise [26, 6].
@@
Furthermore, many state-of-the-art systems default to some variant of the Kinematic Bicycle Model (KBM) [13, 4, 12, 11] due to its simplicity and low computational cost, performing well under low-speed, low-curvature conditions. However, the KBM yields unrealistic trajectories when MMVs perform aggressive or high-curvature maneuvers, such as sharp turns, hard braking, and rapid lane changes. The model does not capture tire slip, load transfer, or rider lean, and its single-track geometry is not generalizable to other forms of micro-mobility with alternate wheel configurations. These limitations make the KBM insufficient for simulating the full range of MMV behaviors that autonomous systems must anticipate in mixed traffic.
@@
  • An interactive evaluation and visualization Framework for GM3 systems. The framework enables plug-and-play specification of MMV dynamics models, vehicle geometries, and parameters (e.g. vehicle, tire, lean) and provides real-time control and trajectory visualization.
@@ II RELATED WORKS
Kinematic Bicycle Model (KBM) is a simple and computationally inexpensive model that is accurate for low speeds [10]. However, it has limitations in modeling high-speed and high-curvature scenarios. [19] demonstrated that KBM only yielded consistent results for lateral acceleration $a_{y}<0.5\mu g$. In an experiment that aimed to improve KBM for higher lateral accelerations  [14], the results show vehicle heading error growing significantly with the curvature due to under-steering. This is because KBM does not account for lean or lateral slip. KBM is also not generalizable to other MMVs with different wheel configurations.
Using a Tire Model as a basis, vehicles with varying numbers of wheels in different positions can be simulated by applying the same tire model to each tire and transforming the resulting forces into the vehicle body frame. [23] validated the brush model with a Chrysler Voyager, and the experimental trajectories closely matched the corresponding simulation results for low speeds. In [17], the brush model and a thermo-physical brush model were validated on a Formula SAE vehicle in extreme scenarios, such as high-speed lane changes. The qualitative and quantitative results show agreement between both models and the experimental data, with an average fitting error of 6.19% for the basic brush model and 6.09% for the thermo-physical model. The semi-empirical Magic Formula – widely used for commercial purposes – was also evaluated against the data and only provided a 0.46% improvement over the brush model. These results demonstrate the viability of the brush model for simulating vehicle dynamics at low and high speeds.
Several implementations of the brush model and other tire models have been designed for specific vehicles. Project Chrono [22] provides a Vehicle class with templates for subsystems, including suspension, steering, and braking, and options for three different classes of tire models: rigid, semi-empirical, and finite element. The user can define a vehicle as a list of axles, each consisting of a set of tires with their own parameters. While this library can handle various wheel configurations, its purpose is to model complex military vehicles. Hence, it does not address rider dynamics, and the computational overhead is excessive for MMVs. There is no general model for MMVs that allows the user to customize the wheel layout similarly to Project Chrono, and this paper intends to fill that gap.
@@ III PRELIMINARIES
The brush model is derived from calculating the deflection of each bristle and integrating over the contact patch between the tire and the road surface, assuming a parabolic pressure distribution and a flat road surface for simplicity. Tables I and II include the minimum parameters and variables required to calculate slip forces for a single tire.
@@
$\ell$ Half contact length m
@@
The brush model uses a 3D coordinate system with $x$ pointing forward along the tire’s trajectory (longitudinal), $y$ pointing right perpendicular to $x$ (lateral), and $z$ pointing up. For simplicity, we assume an isotropic model in which the tire and road properties are the same in all directions ($c_{p}=c_{px}=c_{py}$, where $c_{p}$ is the tread stiffness, and $\mu=\mu_{x}=\mu_{y}$, where $\mu$ is the friction coefficient), and a flat road surface. We define the theoretical slip vector $\mathbf{\sigma}$ with practical slip quantities $s_{x}=-\frac{V_{sx}}{V_{x}}$ and $\tan\alpha=-\frac{V_{sy}}{V_{x}}$ (where $V_{x}$ and $V_{y}$ are the linear longitudinal and lateral velocities, $V_{sx}$ and $V_{sy}$ are the longitudinal and lateral slip velocities, $s_{x}$ is the longitudinal slip ratio, and $\alpha$ is the slip angle).
$\mathbf{\sigma}=-\frac{\mathbf{V_{s}}}{V_{r}}=\begin{pmatrix}\sigma_{x}\\ \sigma_{y}\end{pmatrix}\\ =\begin{pmatrix}\frac{s_{x}}{1+s_{x}}\\ \frac{\tan\alpha}{1+s_{x}}\end{pmatrix}$ (1)
where $\mathbf{V_{s}}$ is the slip velocity vector and $V_{r}=R\Omega$ is the rolling velocity. We introduce spin $\varphi$, which represents the turn slip that occurs when a tire follows a curved path. For a tire with camber angle $\gamma$, radius $R$, turning radius $r_{\text{turn}}$, and reduction factor $\varepsilon_{\gamma}$ (close to 0 for small tires and close to 1 for large tires), the turn slip is
$\varphi=-\frac{1}{r_{\text{turn}}}+(1-\varepsilon_{\gamma})\frac{\sin(\gamma)}{R}.$ (2)
@@
$F_{x}$ $=\begin{cases}\mu F_{z}\frac{\sigma_{x}}{\sigma}[3\psi\sigma-3(\psi\sigma)^{2}+(\psi\sigma)^{3}],&\sigma\leq\frac{1}{\psi}\\ \mu F_{z}\frac{\sigma_{x}}{\sigma},&\sigma>\frac{1}{\psi}\end{cases}$ (4)
$F_{y}$ $=\begin{cases}\mu F_{z}[3\psi^{*}\sigma_{y}-3(\psi^{*}\sigma_{y})^{2}+(\psi^{*}\sigma_{y})^{3}]+\frac{2}{3}c_{p}\ell^{3}\varphi,&|\sigma_{y}|\leq\frac{1}{\psi^{*}}\\ \mu F_{z}\text{sgn}(\alpha),&|\sigma_{y}|>\frac{1}{\psi^{*}}\end{cases}$ (5)
$M_{z}$ $=\begin{cases}-\mu F_{z}\ell\psi^{*}\sigma_{y}[1-3|\psi^{*}\sigma_{y}|+3|\psi^{*}\sigma_{y}|^{2}-|\psi^{*}\sigma_{y}|^{3}],&|\sigma_{y}|\leq\frac{1}{\psi^{*}}\\ 0,&|\sigma_{y}|>\frac{1}{\psi^{*}}\end{cases}$ (6)
The values $\frac{1}{\psi}$ and $\frac{1}{\psi^{*}}$ represent the boundary between the adhesion region ($\sigma\leq\frac{1}{\psi}$), where the bristles deform without sliding, and the sliding region ($\sigma>\frac{1}{\psi}$), where the tire has reached its maximum available friction. Here, sgn is the sign function, $\sigma=\sqrt{\sigma_{x}^{2}+\sigma_{y}^{2}}$, and $\psi^{*}$ is the modified tire parameter for camber angle:
$\psi^{*}=\frac{\psi}{1-\ell\varphi\psi\text{sgn}(\alpha)}$ (7)
@@ IV METHODOLOGY
Figure 2: The GM3 State gets updated with the following steps: (1) load transfer and leaning angle to steer conversion for platforms, (2) control assignment for each tire, (3) individual tire processing with the brush model, and (4) force integration and rider lean force application.
@@ IV-B Load Transfer
$F_{z,front}=\frac{mgl_{r}}{n_{f}L}\qquad F_{z,rear}=\frac{mgl_{f}}{n_{r}L}$ (15)
@@
$T_{long}=\frac{ma_{x}h_{cg}}{L}\qquad T_{lat}=\frac{ma_{y}h_{cg}}{W}$ (16)
@@ IV-D Skateboard Truck Geometry
$\delta_{f}=k\phi\sin{\beta}\qquad\delta_{r}=-k\phi\sin{\beta}$ (21)
@@ IV-E Generalization to Diverse Vehicle Platforms
Our GM3 is widely applicable to model and simulate a diverse range of micro-mobility vehicles across arbitrary layouts. Some common modes are shown in Fig. 6, including bicycle, scooter, skateboard, cart/LSV, unicycle, hoverboard, delta 3-wheeler, tadpole 3-wheeler, 5-wheeled circular platform, and drawn carriage.
Currently, our model supports three control interfaces: explicit manual control, lean-to-steer with truck geometry, and differential drive. Explicit manual control entails a direct connection between the steering input and the wheels, i.e. turning the handlebars or steering wheel causes the wheels to turn by a proportionate angle. This is the most common control interface and is used for bicycle, scooter, cart, delta 3-wheeler, tadpole 3-wheeler, 5-wheeled circular platform, and drawn carriage. Lean-to-steer control couples lean angle to the steering angle of the wheels as described in section IV-D, and is used for skateboard. Lastly, differential drive rotates the wheels at different rates to turn. This control interface is used for hoverboard and can be extended to robots with fixed wheels.
Since GM3 is modular, the control module can be extended to other control interfaces. This allows GM3 to support any wheeled MMV or robot, given individual wheel steering angles and angular velocity from the control module.
((a)) Bicycle (2w)
@@
((f)) Delta (3w)
((g)) Tadpole (3w)
((h)) Circular Platform (5w)
@@
IV-F Interactive Evaluation and Visualization Framework
To evaluate the proposed GM3 and facilitate comparison with the KBM baseline, we developed an interactive, model-agnostic evaluation and visualization framework 7. It is designed to be both a tool for validating the dynamics of different micro-mobility modes and a visualization environment for comparing trajectories produced by different models. Additionally, while this work primarily focuses on comparing GM3 against KBM, the framework is designed to be extensible so additional models can be introduced by adhering to the same interface for control mappings, state, and visual updates.
@@ IV-F Interactive Evaluation and Visualization Framework
At runtime, dynamics are advanced with fixed-step fourth-order Runge-Kutta (RK4) integration. At each time step, inputs or controls are read (from keyboard or a script), transformed into model-specific commands, and fed to the appropriate model to update the state. The new state is then used to append time-aligned state and control values to logs and update the visualization which, beyond vehicle effects, extends an accumulating trajectory trace. The trace can be cleared or reset to rerun maneuvers with different configurations. Run histories can be exported to CSV with model-specific fields for analysis.
Figure 7: Micro-mobility evaluation and visualization framework diagram.
@@ V EXPERIMENTS
In our experiments, we evaluate the accuracy and correctness of GM3 in reproducing realistic, physically-ground micro-mobility trajectories, as compared to the KBM. Additionally, we stress test both models across different aggressive maneuvers, including S-curves, U-turns, and left & right turns.
@@
Dataset. Experiments are conducted using the Stanford Drone Dataset (SDD) [20], a large collection of aerial videos recorded over the Stanford campus. Each video is accompanied by frame-level annotations in the form of axis-aligned bounding boxes and categorical labels for eight classes of road users, including pedestrians, cyclists, skaters, carts, cars, and buses. For this work specifically, we focus on the deathCircle (roundabout) scene since it contains diverse micro-mobility interactions as bicyclists, skateboarders, and carts navigate a shared space alongside pedestrians taking high curvature routes. Image-plane trajectories are converted to metric world coordinates using the scene’s pixel-to-meter scale factors [2]. For each annotation, we extract the bottom-midpoint of the bounding box, then resample at the native 30 fps [3] to obtain a uniform step size $\Delta t\approx 0.03$s.
Additionally, to demonstrate GM3’s performance on high-speed scenarios, we generate a set of trajectories with the following maneuvers: S-curve, left/right turn, and left/right U-turn. The resulting positions were computed using a model-agnostic dynamics model rather than directly integrating the full GM3 or KBM equations. The state is $(x,y,\theta,v)$, with commanded acceleration $a_{\mathrm{cmd}}(t)$ and curvature $\kappa_{\mathrm{cmd}}(t)$ passed through first-order dynamics, $\dot{a}=(a_{\mathrm{cmd}}-a)/\tau_{a}$ and $\dot{\kappa}=(\kappa_{\mathrm{cmd}}-\kappa)/\tau_{\kappa}$, to produce smooth control response. The constants $\tau_{a}$ and $\tau_{\kappa}$ are response times for acceleration and curvature, respectively. Vehicle speed evolves as $\dot{v}=a-dv$ (with linear drag), heading as $\dot{\theta}=v\kappa$, and position as $\dot{x}=v\cos\theta$, $\dot{y}=v\sin\theta$. To remain physically realistic, curvature is limited by tire-road friction via $|\kappa|\leq\mu g/v^{2}$, so aggressive commands are clipped to feasible lateral acceleration.
@@
Vehicle Instantiation. Vehicles, one per mode, are instantiated using representative, publicly reported parameters (e.g., wheelbase, track width, tire radius, mass) taken from widely available bicycle, skateboard, and cart specifications on company websites. The bicycle MMV agent uses a single track geometry with lean, the skater uses a skateboard configuration with rider lean, and the cart uses a four-wheel layout and no rider lean. All layouts use load transfer. To prevent overfitting, parameters are fixed per mode and not tuned per track.
@@
Metrics. For each trajectory, we quantify model performance with two measures: (1) Average Displacement Error (ADE) and (2) Discrete Fréchet Distance (DFD). Average Displacement Error (ADE) is computed as the mean Euclidean Distance between the estimated model positions and the corresponding ground truth positions across all time steps. ADE captures how well the simulated trajectory remains aligned with the true path when both are driven by the same underlying control sequence. To account for overall path similarity, we report discrete Fréchet distance (DFD).
@@
Results. Table III summarizes ADE and DFD errors across all three modes. GM3 reduces ADE relative to KBM across all modes (Biker: -5.16%, Skater: -8.77%, Cart: -6.82%), which indicates better step-by-step alignment. For DFD, GM3 improves on the Skater (-6.1%) and Cart (-2.34%) modes, but is worse on Biker (+11.56%). Qualitatively, GM3 reproduces better short-term responses to controls due to the tire-level formulation, which aligns with the ADE observations. Since Fréchet distance emphasizes global path similarity, the larger lean angles of cyclists relative to skaters and golf carts introduce lateral biases that accumulate along the extracted bottom-midpoint trajectories, which can inflate the distance. However, aggregated across all modes, GM3 performs only around 1% worse than KBM with respect to DFD. Furthermore, the campus setting primarily captures low-speed, cooperative interactions on narrow walkways and roadways (even at the roundabout), so hard braking, rapid lane changes, and sustained high-speed, high-curvature turns tend to be rare. As a result, the reported differences likely understate the advantage GM3 provides in higher-speed scenarios with larger slip angles (e.g., bike lanes, downhill stretches, large evasive maneuvers). Please refer to the supplementary video for live side-by-side GM3 and KBM comparisons.
Results in table IV demonstrate that GM3 more accurately reproduces ground truth aggressive micro-mobility maneuvers than KBM. GM3 achieves lower ADE and DFD errors across the five maneuver classes with the biggest improvements observed for aggressive left and right turns. This behavior is also reflected in 8, where we see GM3 closely follows the ground-truth paths while KBM tends to overshoot or understeer in high-speed and high-curvature segments. These observations can primarily be explained by GM3’s modeling of slip forces and load transfer, which allow it to better capture tire-level dynamics.
@@
[…]
Maneuver GM3 KBM GM3 KBM
S-curve 0.7398 2.7537 4.0851 4.4777
Turn (Left) 0.2135 5.0892 0.3995 15.5169
Turn (Right) 0.2786 5.0892 0.4982 15.5169
U-turn (Left) 2.9065 6.7858 5.1678 26.0783
U-turn (Right) 2.8124 6.7858 4.8561 26.0783
TABLE IV: Mean Average Displacement Error (ADE) and discrete Fréchet distance (DFD) across 5 aggressive maneuver classes.
((a)) S-Curve
((b)) U-turn
((c)) Right turn
Figure 8: GM3 and KBM trajectory reconstruction for S-curve, U-turn, and Right turn maneuvers, compared to the captured ground truth trajectories for these scenarios.